OpenStudy (anonymous):

simplify. assume that the expression under the radical represents a nonnegative number. sqrt(64x^2)

OpenStudy (anonymous):

\[8x\]

OpenStudy (anonymous):

since \[(8x)^2=8^2x^2=64x^2\]

OpenStudy (anonymous):

8x as satellite has said^^^^^^^^^^^^^^

OpenStudy (anonymous):

8x